OpenStudy (anonymous):

simplify radical 5 (10- 4 radical 2)

OpenStudy (anonymous):

can you try to be a little more descriptive or write it out for me

OpenStudy (anonymous):

\[\sqrt{5}(10-4\sqrt{2}\]

OpenStudy (anonymous):

)

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

aight one sec

OpenStudy (anonymous):

im not quite sure how else to simplify that other than 10*radical5-4*radical10

OpenStudy (anonymous):

do you think you could help me out one more time @jim_thompson5910

jimthompson5910 (jim_thompson5910):

PeachRings is correct. The only thing you can do is distribute to get \[\Large 10\sqrt{5} - 4\sqrt{10}\]

jimthompson5910 (jim_thompson5910):

You can't simplify the radicals further or combine like terms. So you just leave it like that.

OpenStudy (anonymous):

never mind I see my mistake, Thank you @PeachRings & @jim_thompson5910